# “Hello, is this Anna?": A First Look at Pig-Butchering Scams

### v1

@@ Abstract
Pig-butchering scams, or Sha Zhu Pan, have emerged as a complex form of cyber-enabled financial fraud that combines elements of romance, investment fraud, and advanced social engineering tactics to systematically exploit victims. In this paper, we present the first qualitative analysis of pig-butchering scams, informed by in-depth semi-structured interviews with $N=26$ victims. We capture nuanced, first-hand accounts from victims across multiple regions, providing insight into the lifecycle of pig-butchering scams and the complex emotional and financial manipulation involved. We systematically analyze each phase of the scam, revealing that perpetrators employ tactics such as staged trust-building, fraudulent financial platforms, fabricated investment returns, and repeated high-pressure tactics, all designed to exploit victims’ trust and financial resources over extended periods. Our findings reveal an organized scam lifecycle characterized by emotional manipulation, staged financial exploitation, and persistent re-engagement efforts that amplify victim losses. We also find complex psychological and financial impacts on victims, including heightened vulnerability to secondary scams. Finally, we propose actionable intervention points for social media and financial platforms to curb the prevalence of these scams and highlight the need for non-stigmatizing terminology to encourage victims to report and seek assistance.
@@ 1 Introduction
  • We present a comprehensive anatomy of pig-butchering scams, breaking down the lifecycle of the scam into distinct stages, from initial contact through financial exploitation, and potentially onto an extended “encore" scam.
  • Through in-depth interviews with scam victims, we provide a detailed understanding of the emotional and psychological tolls, including shame, trust issues, and mental health impacts, offering insights into the broader consequences of these scams beyond financial loss.
@@ 2.2 Pig-Butchering Scams
Pig-butchering scams, also known as Sha Zhu Pan, are a form of online fraud where perpetrators cultivate long-term relationships with victims to defraud them financially [10, 31]. This method relies heavily on social engineering techniques, where trust is cultivated over time, making victims more susceptible to manipulation [25]. Prior work [26] has highlighted the psychological aspects of these scams, noting that victims often exhibit low self-esteem, particularly in interpersonal-social dimensions, which scammers exploit to establish emotional trust before executing the fraud. Complementing this work, Wang and Zhou [46] provide insights into the scammers’ persuasive techniques, revealing that these scams follow a premeditated, structured approach to manipulate victims emotionally and financially. Legal responses to pig-butchering scams have been varied, particularly in regions like China, where the phenomenon has gained notoriety [45]. From a legal standpoint, Tan [44] has found that the difficulty in identifying accomplices and collecting evidence complicates the prosecution of these cybercrimes, limiting the effectiveness of law enforcement. The Chinese government has initiated campaigns aimed at curbing these scams, yet the problem persists and has expanded beyond national borders, affecting victims in Southeast Asia, North America, and beyond.
@@ 2.3 Users & Scams
Studying scams through the lens of user experience provides valuable insights into the scam anatomy as well as effective prevention and intervention strategies. Miramirkhani et al [33] conducted a large-scale analysis of technical support scams, which revealed that these scams exploit users’ lack of technical knowledge, often through malvertising and fake tech support websites. Chen et al [13] explored the factors leading to scam victimization, finding that routine internet activities and low self-control significantly increase the risk of being scammed. Research on online romance scams [43, 14] shows that victims often experience a double trauma of financial loss and emotional betrayal, which can lead to long-term psychological harm. Understanding security experiences and behaviors of at-risk user groups has also been a focal point of research, showing how certain unique contextual factors amplify the risk of cybercrime for certain demographics [6, 47, 20, 4]. Herbert et al [24] compare and contrast cybercrime experiences between various at-risk user groups. Qualitative research has proven to be valuable tools in S&P research [37, 42, 40, 17]. The unique nature of pig butchering scams -— combining elements of romance, investment fraud, and advanced cyber tactics necessitates a focused analysis to understand its lifecycle fully. To this end, we aim to extend prior qualitative research on online scams to discover the specific strategies, psychological manipulation, and long-term grooming tactics that define pig-butchering.
@@ 4.1 The Lure
Cold Texts. A significant proportion of participants ($n=18$, 69.2%) reported that the initial contact with the scammer was through unsolicited messages via SMS, WhatsApp, or other messaging applications, where the scammer pretended to have entered an incorrect number. The first message of the form “Hello, is this X", where X was a random name. After being corrected by the victim, they would apologize but still introduce themselves and ask to continue the conversation.
Dating Apps. 6 participants (23%) were approached on dating platforms (2 via Tinder, 3 via Bumble, and 1 via Hinge). Scammers created attractive profiles, often featuring professionally taken photographs and bios highlighting shared interests with potential victims. Once a match was made, the scammers initiated conversations that quickly became personal, and attempted to move them off-platform on a messaging service like WhatsApp or Telegram.
Social Media. A smaller segment ($n=2$) encountered scammers through social media platforms (1 each via Instagram and LinkedIn). Scammers would send connection requests or follow the victims, often engaging by commenting on public posts, and initiate conversation via personal messages. Scammer profiles were carefully crafted and not obviously fake; P4 recounts:
@@ 4.2 The Bond
Nature of Relationships. Scammers cultivated either business or romantic relationships with victims. The majority of participants ($n=16$, 61.5%) developed what they believed to be romantic relationships with the scammers. These relationships were characterized by frequent communication, expressions of affection, and discussions about future plans together. The remaining 10 participants (38.5%) participants were enticed through the prospect of business partnerships or mentorships. Scammers presented themselves as successful entrepreneurs or seasoned professionals willing to share their knowledge.
Rapport Building Techniques. Scammers invested significant time in building rapport, with participants noting daily communication over several weeks or months. They shared personal stories ($n=8$, 30.7%), discussed family and career aspirations ($n=11$, 42.3%) and showed empathy towards the victims’ experiences ($n=5$, 19.2%). Scammers frequently used mirroring as a manipulation technique; 9 participants (34.6%) noted that the scammer agreed with them, or reacted positively to any view of theirs. In the words of P8:
@@
Scammers also gradually disclosed personal information to build trust, sometimes sharing fabricated hardships or vulnerabilities, such as a challenging past ($n=4$, 15.38%) or experiences with debt ($n=9$, 34.6%).
Trust Building Techniques. Some scammers went even a step further in order to build more trust with victims. 17 participants (65.38%) reported that the scammer sent frequent pictures of themselves, and their surroundings. 7 participants (26.9%) reported having audio calls, and 3 participants (11.5%) video calls with the scammers; they were able to verify that the pictures sent matched with the person they spoke to on the video call. According to P2:
@@
Social Media. Scammers used well-designed social media profiles to keep up the appearance of a genuine personality, and to back up some of the claims they made to victims. 20 participants (77%) reported having connected with the scammer on at least one other social media platform: Facebook ($n=10$, 38.46%), Instagram ($n=14$,53.84%), or X ($n=9$, 34.61%). Participants reported that the profiles were consistent with the details that the scammer shared, and that this was an important factor in them believing that the scammer was in fact a genuine person. For example, P18 shared:
@@
Duration. The bond phase showed that scammers had tremendous patience; this phase lasted anywhere from 3 to 11 months before the scammer moved on to the next stage of the scam. Scammers spent a significant amount of time bonding with the victim, to cultivate a sense of trust. According to 20 participants (76.9%), this led them to letting their guard down and dispel any suspicions about this being a scam; P5 commented:
@@ 4.3 The Bait
Introducing Investments. Scammers typically introduce cryptocurrency investment into the conversation in a casual manner, referring to it as an extra source of income ($n=15$, 57.7%) or even a hobby ($n=3$, 11.5%). They may claim to have insider knowledge or expertise in crypto trading ($n=19$, 73%), or an investment consultant who manages their portfolio ($n=7$, 27%). The bait is framed as an exclusive opportunity, often portrayed as an investment that is only available to a select few. 6 participants reported that the scammer described their investments in a manner aligned with the victims’ circumstances, such as debts or future aspirations. In the words of P23:
@@
Providing Social Proof. Scammers provided fake screenshots or access to counterfeit investment platforms displaying impressive returns. 22 participants (84.6%) were shown evidence of significant profits through fabricated screenshots. Scammers may also provide proof via testimonials; 5 participants (19.2%) reported that the scammer showed them emails or message screenshots confirming successful returns on investments. Further, 4 participants (15.4%) were also given emails and phone numbers of other supposed investors, so that they could act as references. Participant P20 says:
@@
Leveraging FOMO. The scammer emphasizes the exclusivity and limited nature of the opportunity, thereby exploiting the victim’s fear of missing out (FOMO). This approach is designed to make the victim feel privileged to be “invited" into the scheme, which heightens the perceived legitimacy and desirability of the investment; 18 participants (69.2%) reported that the scammer presented the investment guidance as secret or highly specialized. At this point, the victim, already emotionally invested and trusting of the scammer, is primed to believe that this opportunity is both genuine and lucrative. {mdframed}[style=insightstyle] After a bond has been established, the scammer introduces a fraudulent investment opportunity, typically framed as a low-risk, high-reward endeavor, such as cryptocurrency trading. The scammer uses fabricated success stories and social proof to persuade the victim of the legitimacy of the offer. This phase is designed to exploit the trust and rapport built in earlier stages, positioning the scammer as a trusted financial advisor or insider.
@@ 4.4 The Feed
Fraudulent Investment Platforms. Scammers provide access to sophisticated-looking but fraudulent investment platforms or applications, and asked victims to sign up for an account. All the participants reported creating accounts on such websites. 11 participants (42.3%) reported that the scammer helped them set up their accounts through step-by-step instructions; 5 (19.2%) reported that there were even help articles or some form of customer support. These platforms were authentic-looking, and displayed real-time information, with efforts taken to ensure that they resembled real websites. P22 says:
@@
Initial Investments. The scammer typically begins by encouraging the victim to make a modest initial investment. Participants reported initial investments of \$20 – \$125. This allowed the victim to test the waters without feeling at great risk, and it also provides an opportunity for the scammer to build credibility. To maintain the illusion of legitimacy, the scammer may manipulate fake trading platforms or falsify investment returns, making it appear as if the initial investment has yielded positive gains; all the participants saw significant gains in their initial investment (2x – 5x). 3 participants (11.5%) also reported being allowed to withdraw some amount, further convincing them of the credibility of the investments. As P1 says:
@@
Escalating Investments. Scammers applied subtle pressure to reinvest profits rather than withdraw them. 23 participants (88.4%) were encouraged to roll over their gains into new investments, so as to earn even more. As victims grew more comfortable, scammers suggested larger investments to achieve greater profits. According to 13 participants (50%), victims were encouraged to use their ’profits’ as leverage for higher-tier investment opportunities. All participants reported increasing their investment amounts significantly (\$3,500 – \$65,00) by putting in more money. Participants also saw tremendous gains on their investments on the platform (2x – 12x).
Manipulation Techniques. Scammers maintained regular communication, celebrating the victim’s ’successes’ and reinforcing the notion of smart investing; 12 participants (46.1%) received congratulatory messages whenever their account showed any gains. Scammers often claimed to be investing alongside the victims, fostering a sense of shared endeavor; they manipulated the investment platform so as to create this illusion, even ’lending’ money if needed ($n=14$, 53.8%). P14 recounts:
@@
Further, according to 8 participants (30.7%), scammers introduced complex investment strategies, using technical jargon to appear knowledgeable. Finally, all participants reported that scammers discouraged them from discussing investments with others, citing reasons like market confidentiality. {mdframed}[style=insightstyle] Scammers persuade victims to invest in fraudulent platforms that mimic legitimate financial services. Initial investments are typically small, with fabricated returns used to encourage further contributions. The scammer maintains ongoing communication, reinforcing the illusion of success and increasing the victim’s financial commitment by encouraging reinvestment of perceived gains.
@@ 4.5 The Squeeze
Exclusive Deals. After continuing investments, all participants were presented with time-sensitive, high-return investment opportunities. Scammers claimed access to insider deals ($n=12$, 46.2%) or initial coin offerings (ICOs) ($n=9$, 34.6%) about to explode in value. These investments had minimum amounts, which were often large (in the tens of thousands). 18 participants(69.2%) reported that they felt inclined to invest, because of the perceived returns so far.
Pressure Tactics. To ensure the high investment, scammers increased the frequency and urgency of their messages. All participants reported receiving multiple messages or calls daily, emphasizing the need for immediate action. Scammers employed emotional manipulation, expressing disappointment or suggesting that the victim was missing out due to hesitation. 11 participants (42.3%, for whom the relationship was of a romantic nature) also reported frustruation on the scammer’s part; as P10 put it:
@@
Continuing Investments. Scammers encouraged victims to “re-invest" their gains, and add more funds rather than attempt to withdraw them. Exclusive opportunities as described above keep on occurring till the victim invests. The squeeze serves to drain any remaining financial resources the victim may have access to, pushing them to borrow money or liquidate other assets if necessary. 9 participants (34.6%) reported that the scammer recommended using credit cards or applying for new credit cards so as to use credit to invest. The squeeze phase lasts for as long as the victim is willing to invest; participants reported continuing large investments and perceived profit for 1-7 months. By the end of this phase, the scammer aims to have extracted as much financial value as possible.

### v2

# “Hello, is this Anna?": Unpacking the Lifecycle of Pig-Butchering Scams

@@ Abstract
Pig-butchering scams have emerged as a complex form of fraud that combines elements of romance, investment fraud, and advanced social engineering tactics to systematically exploit victims. In this paper, we present the first qualitative analysis of pig-butchering scams, informed by in-depth semi-structured interviews with $N=26$ victims. We capture nuanced, first-hand accounts from victims, providing insight into the lifecycle of pig-butchering scams and the complex emotional and financial manipulation involved. We systematically analyze each phase of the scam, revealing that perpetrators employ tactics such as staged trust-building, fraudulent financial platforms, fabricated investment returns, and repeated high-pressure tactics, all designed to exploit victims’ trust and financial resources over extended periods. Our findings reveal an organized scam lifecycle characterized by emotional manipulation, staged financial exploitation, and persistent re-engagement efforts that amplify victim losses. We also find complex psychological and financial impacts on victims, including heightened vulnerability to secondary scams. Finally, we propose actionable intervention points for social media and financial platforms to curb the prevalence of these scams and highlight the need for non-stigmatizing terminology to encourage victims to report and seek assistance.
Copyright is held by the author/owner. Permission to make digital or hard copies of all or part of this work for personal or classroom use is granted without fee.
USENIX Symposium on Usable Privacy and Security (SOUPS) 2025.
August 10–12, 2025, Seattle, WA, United States.
@@ 1 Introduction
  • We present a comprehensive anatomy of pig-butchering scams, breaking down the lifecycle of the scam into seven distinct stages, from initial contact through financial exploitation, and potentially onto an extended “encore" scam.
  • Through in-depth interviews with scam victims, we provide a detailed understanding of the emotional and psychological tolls, including shame, trust issues, and mental health impacts, offering qualitative insights into the broader consequences of these scams beyond financial loss.
@@ 2.2 Pig-Butchering Scams
Pig-butchering scams, also known as Sha Zhu Pan, are a form of online fraud where perpetrators cultivate long-term relationships with victims to defraud them financially [10, 31]. This method relies heavily on social engineering techniques, where trust is cultivated over time, making victims more susceptible to manipulation [25]. Prior work [26] has highlighted the psychological aspects of these scams, noting that victims often exhibit low self-esteem, particularly in interpersonal-social dimensions, which scammers exploit to establish emotional trust before executing the fraud. Complementing this work, Wang and Zhou [46] provide insights into the scammers’ persuasive techniques. While they offer a valuable conceptual framing of pig-butchering scams through a three-stage model, our work substantially expands this foundation through empirical depth and granularity. They rely on publicly available reports to outline scammer strategies, whereas our study draws from 26 in-depth, first-hand interviews with scam victims, capturing the evolving tactics of perpetrators in real-world cases which go beyond their model. Legal responses to pig-butchering scams have been varied, particularly in regions like China, where the phenomenon has gained notoriety [45]. From a legal standpoint, Tan [44] has found that the difficulty in identifying accomplices and collecting evidence complicates the prosecution of these cybercrimes, limiting the effectiveness of law enforcement.
@@ 2.3 Users & Scams
Studying scams through the lens of user experience provides valuable insights into the scam anatomy as well as effective prevention and intervention strategies. Miramirkhani et al [33] conducted a large-scale analysis of technical support scams, which revealed that these scams exploit users’ lack of technical knowledge, often through malvertising and fake tech support websites. Chen et al [13] explored the factors leading to scam victimization, finding that routine internet activities and low self-control significantly increase the risk of being scammed. Research on online romance scams [43, 14] shows that victims often experience a double trauma of financial loss and emotional betrayal, which can lead to long-term psychological harm. Understanding security experiences and behaviors of at-risk user groups has also been a focal point of research, showing how certain unique contextual factors amplify the risk of cybercrime for certain demographics [6, 47, 20, 4]. Herbert et al [24] compare and contrast cybercrime experiences between various at-risk user groups. Qualitative research has proven to be valuable tools in S&P research [37, 42, 40, 17]. The unique nature of pig butchering scams combining elements of romance, investment fraud, and advanced cyber tactics necessitates a focused analysis to understand its lifecycle fully. To this end, we aim to extend prior qualitative research on online scams to discover the specific strategies, psychological manipulation, and long-term grooming tactics that define pig-butchering.
@@ 4.1 The Lure
Cold Texts. A significant proportion of participants reported that the initial contact with the scammer was through unsolicited messages via SMS, WhatsApp, or other messaging applications, where the scammer pretended to have entered an incorrect number. The first message of the form “Hello, is this X", where X was a random name. After being corrected by the victim, they would apologize but still introduce themselves and ask to continue the conversation.
Dating Apps. 6 participants were approached on dating platforms (Tinder, Bumble, and Hinge). Scammers created attractive profiles, often featuring professionally taken photographs and bios highlighting shared interests with potential victims. Once a match was made, the scammers initiated conversations that quickly became personal, and attempted to move them off-platform on a messaging service like WhatsApp or Telegram.
Social Media. A smaller segment encountered scammers through social media platforms (Instagram and LinkedIn). Scammers would send connection requests or follow the victims, often engaging by commenting on public posts, and initiate conversation via personal messages. Scammer profiles were carefully crafted and not obviously fake; P4 recounts:
@@ 4.2 The Bond
Nature of Relationships. Scammers cultivated either business or romantic relationships with victims. The majority of participants developed what they believed to be romantic relationships with the scammers. These relationships were characterized by frequent communication, expressions of affection, and discussions about future plans together. The remaining participants were enticed through the prospect of business partnerships or mentorship. Scammers presented themselves as successful entrepreneurs or seasoned professionals willing to share their knowledge.
Rapport Building Techniques. Scammers invested significant time in building rapport, with participants noting daily communication over several weeks or months. They shared personal stories, discussed family and career aspirations and showed empathy towards the victims’ experiences. Scammers frequently used mirroring as a manipulation technique; participants noted that the scammer agreed with them, or reacted positively to any view of theirs. In the words of P8:
@@
Scammers also gradually disclosed personal information to build trust, sometimes sharing fabricated hardships or vulnerabilities, such as a challenging past or experiences with debt.
Trust Building Techniques. Some scammers went even a step further in order to build more trust with victims. Participants reported that the scammer sent frequent pictures of themselves, and their surroundings. 7 participants reported having audio calls, and 3 participants video calls with the scammers; they were able to verify that the pictures sent matched with the person they spoke to on the video call. According to P2:
@@
Social Media. Scammers used well-designed social media profiles to keep up the appearance of a genuine personality, and to back up some of the claims they made to victims. Most participants (77%) reported having connected with the scammer on at least one other social media platform: Facebook, Instagram, or X. Participants reported that the profiles were consistent with the details that the scammer shared, and that this was an important factor in them believing that the scammer was in fact a genuine person. For example, P18 shared:
@@
Duration. The bond phase showed that scammers had tremendous patience; this phase lasted anywhere from 3 to 11 months before the scammer moved on to the next stage of the scam. Scammers spent a significant amount of time bonding with the victim, to cultivate a sense of trust. According to most participants, this led them to letting their guard down and dispel any suspicions about this being a scam; P5 commented:
@@ 4.3 The Bait
Introducing Investments. Scammers typically introduce cryptocurrency investment into the conversation in a casual manner, referring to it as an extra source of income or even a hobby. They may claim to have insider knowledge or expertise in crypto trading, or an investment consultant who manages their portfolio. The bait is framed as an exclusive opportunity, often portrayed as an investment that is only available to a select few. 6 participants reported that the scammer described their investments in a manner aligned with the victims’ circumstances, such as debts or future aspirations. In the words of P23:
@@
Providing Social Proof. Scammers provided fake screenshots or access to counterfeit investment platforms displaying impressive returns. Most participants were shown evidence of significant profits through fabricated screenshots. Scammers may also provide proof via testimonials; 5 participants reported that the scammer showed them emails or message screenshots confirming successful returns on investments. Further, some were also given emails and phone numbers of other supposed investors, so that they could act as references. Participant P20 says:
@@
Leveraging FOMO. The scammer emphasizes the exclusivity and limited nature of the opportunity, thereby exploiting the victim’s fear of missing out (FOMO). This approach is designed to make the victim feel privileged to be “invited" into the scheme, which heightens the perceived legitimacy and desirability of the investment; most participants reported that the scammer presented the investment guidance as secret or highly specialized. At this point, the victim, already emotionally invested and trusting of the scammer, is primed to believe that this opportunity is both genuine and lucrative. {mdframed}[style=insightstyle] After a bond has been established, the scammer introduces a fraudulent investment opportunity, typically framed as a low-risk, high-reward endeavor, such as cryptocurrency trading. The scammer uses fabricated success stories and social proof to persuade the victim of the legitimacy of the offer. This phase is designed to exploit the trust and rapport built in earlier stages, positioning the scammer as a trusted financial advisor or insider.
@@ 4.4 The Feed
Fraudulent Investment Platforms. Scammers provide access to sophisticated-looking but fraudulent investment platforms or applications, and asked victims to sign up for an account. All the participants reported creating accounts on such websites. 11 participants reported that the scammer helped them set up their accounts through step-by-step instructions; a few reported that there were even help articles or some form of customer support. These platforms were authentic-looking, and displayed real-time information, with efforts taken to ensure that they resembled real websites. P22 says:
@@
Initial Investments. The scammer typically begins by encouraging the victim to make a modest initial investment. Participants reported initial investments of \$20 – \$125. This allowed the victim to test the waters without feeling at great risk, and it also provides an opportunity for the scammer to build credibility. To maintain the illusion of legitimacy, the scammer may manipulate fake trading platforms or falsify investment returns, making it appear as if the initial investment has yielded positive gains; all the participants saw significant gains in their initial investment (2x – 5x). 3 participants also reported being allowed to withdraw some amount, further convincing them of the credibility of the investments. As P1 says:
@@
Escalating Investments. Scammers applied subtle pressure to reinvest profits rather than withdraw them. Most participants were encouraged to roll over their gains into new investments, so as to earn even more. As victims grew more comfortable, scammers suggested larger investments to achieve greater profits. According to 13 participants, victims were encouraged to use their ’profits’ as leverage for higher-tier investment opportunities. All participants reported increasing their investment amounts significantly (\$3,500 – \$65,00) by putting in more money. Participants also saw tremendous gains on their investments on the platform (2x – 12x).
Manipulation Techniques. Scammers maintained regular communication, celebrating the victim’s ’successes’ and reinforcing the notion of smart investing; participants received congratulatory messages whenever their account showed any gains. Scammers often claimed to be investing alongside the victims, fostering a sense of shared endeavor; they manipulated the investment platform so as to create this illusion, even ’lending’ money if needed. P14 recounts:
@@
Further, according to some participants, scammers introduced complex investment strategies, using technical jargon to appear knowledgeable. Finally, all participants reported that scammers discouraged them from discussing investments with others, citing reasons like market confidentiality. {mdframed}[style=insightstyle] Scammers persuade victims to invest in fraudulent platforms that mimic legitimate financial services. Initial investments are typically small, with fabricated returns used to encourage further contributions. The scammer maintains ongoing communication, reinforcing the illusion of success and increasing the victim’s financial commitment by encouraging reinvestment of perceived gains.
@@ 4.5 The Squeeze
Exclusive Deals. After continuing investments, all participants were presented with time-sensitive, high-return investment opportunities. Scammers claimed access to insider deals or initial coin offerings (ICOs) about to explode in value. These investments had minimum amounts, which were often large (in the tens of thousands). Most participants reported that they felt inclined to invest, because of the perceived returns so far.
Pressure Tactics. To ensure the high investment, scammers increased the frequency and urgency of their messages. All participants reported receiving multiple messages or calls daily, emphasizing the need for immediate action. Scammers employed emotional manipulation, expressing disappointment or suggesting that the victim was missing out due to hesitation. Some participants, for whom the relationship was of a romantic nature) also reported frustruation on the scammer’s part; as P10 put it:
@@
Continuing Investments. Scammers encouraged victims to “re-invest" their gains, and add more funds rather than attempt to withdraw them. Exclusive opportunities as described above keep on occurring till the victim invests. The squeeze serves to drain any remaining financial resources the victim may have access to, pushing them to borrow money or liquidate other assets if necessary. 9 participants reported that the scammer recommended using credit cards or applying for new credit cards so as to use credit to invest. The squeeze phase lasts for as long as the victim is willing to invest; participants reported continuing large investments and perceived profit for 1-7 months. By the end of this phase, the scammer aims to have extracted as much financial value as possible.